OpenStudy (anonymous):

reward and medal 5.) a cash box contains 12.25 in quarters, dimes, and nickels. If there were five more than twice as many dimes as nickels and one less than three times as many quarters as nickels, how many of each kind of coin was there? I need two equations.. no answer just two equations. 6.) Vera's piggybank contained 9.35 in quarters, dimes, and nickels. there were six more than two times as many nickels as quarters and four less dimes than quarters. how many of each kind of con was there in the bank?????? I don't need a answer. I need two equations..

OpenStudy (anonymous):

7.) rosa, who works at a fast food restaurant, received 9.05 in tips one afternoon, all in quarters, dimes, and nickels. there were ten less dimes than quarters and five less nickels than dimes. how many of each kind of coin was there?

OpenStudy (anonymous):

I NEED THREE equations per question.. is all

OpenStudy (anonymous):

OKAYYY. Well ill show you how to make the equations for 5 and you should be able to apply the same method to the other ones. Let \(q\) represent quarters Let \(d\) represent dimes Let \(n\) represent nickels So you have \(12.25=q+d+n\) And you have Equation 1. \(2n+5=d\) Equation 2.\(3n-1=q\) So you have \(12.25=(3n-1)+(2n+5)+n\) And you can use this to find nickels and when you do, plug into Equations one and two to find dimes and quarters